OpenStudy (anonymous):

please help [10.02] A map has a scale of 5 km to 3 inches. Which ratio of measures forms a proportion with the map's scale of km to inches? 9 over 7. 17 over 15. 30 over 18. 13 over 11.

OpenStudy (alexandervonhumboldt2):

9/7=5/3 false 17/15=5/3 false

OpenStudy (alexandervonhumboldt2):

30/18=5/3 TRUE

OpenStudy (alexandervonhumboldt2):

13/11=5/3 false

OpenStudy (anonymous):

ok i get how you got that thanks`